# UFGraphFR: Graph Federation Recommendation System based on User Text description features

### v3

@@ Abstract
Federated learning has become a key paradigm that utilizes the "data available but invisible" feature to protect user privacy, especially in the artificial intelligence field where user privacy is highly valued. Federated learning enables users to collaboratively train models without sharing raw data. Motivated by this, federated recommendation systems offer a promising architecture that balances user privacy with recommendation accuracy through distributed collaborative learning. However, existing federated recommendation methods often neglect the underlying semantic or behavioral relationships between users during parameter aggregation, which limits their recommendation effectiveness. To overcome this limitation, graph-based federated recommendation systems have been proposed to leverage neighborhood information. Yet, conventional graph construction methods usually require access to raw user data or explicit social links, which contradicts the strict privacy requirements of federated learning. In this paper, we propose UFGraphFR, a novel model designed to effectively capture user relationships for constructing user graphs while preserving the confidentiality of users’ raw data. (1) To better utilize users’ textual feature descriptions under the hypothesis that users with similar textual profiles exhibit analogous preferences, we propose a Joint-Embedding Layer architecture designed to encode structured user features into fixed-dimensional vector representations. (2) To prevent raw data exposure during temporal pattern extraction, we introduce a Transformer-based sequence encoder with differential privacy guarantees, which employs multi-head self-attention mechanisms to model long-range dependencies in user behavioral sequences through on-device processing. These embeddings are then used to construct a semantic user graph for server-side aggregation via a lightweight graph neural architecture. UFGraphFR ensures privacy-preserving, personalized recommendations while maintaining model robustness and convergence efficiency under heterogeneous conditions. The code is available at: https://github.com/trueWangSyutung/UFGraphFR.
@@
Federated Learning;Federated Recommendation System;Recommendation
journal: Arxiv
[…]
[label1]organization=School of Computer Science and Engineering, addressline=Tianjin University of Technology, city=Tianjin, postcode=300384, state=Tianjin, country=China
{highlights}
Proposed a joint-embedding layer that converts structured user data into privacy-preserving text-based representations.
Developed a differentially private semantic graph construction method to model user correlations in federated learning.
Demonstrated superior performance over both federated and centralized baselines across multiple real-world datasets.
Validated framework effectiveness with different pre-trained language model embeddings.
@@ 1 Introduction
In the era of big data and privacy computing, intelligent recommendation systems have become indispensable infrastructure for modern digital services, enabling personalized movie suggestions, targeted food delivery recommendations, and even emotion-aware content curation through user comment analysis[1]. However, the conventional centralized paradigm that aggregates raw user data on servers poses critical privacy risks, as evidenced by growing user reluctance to share personal information with service providers[2]. In recent years, with the proposal of federated learning, which is a privacy-preserving framework that aims to better protect users’ privacy[3]. More and more research directions have started to apply this framework. Its emergence has catalyzed its application in the field of recommender systems, allowing the training of complete recommender system models without the raw data leaving the client device locally[4] .
Despite significant progress, there are two fundamental limitations of existing joint recommendation methods: (1). They model users as isolated entities mainly through local training goals, ignoring the inherent social correlations and behavioral dependencies among user groups[5], such as belonging to the same Beijing area, belonging to the same male gender, etc. This oversimplification contradicts real-world observations, where users with common interests often exhibit collaborative consumption patterns. (2). Despite the remarkable success of graph-enhanced recommender systems in centralized environments through neighborhood information aggregation, their direct application in federated learning is still prohibited due to privacy violations resulting from explicit user relationship disclosure[6] Combining the problems of the above two points, a graph augmentation based federated recommender system is proposed [7, 8]. (3). Current approaches mainly rely on Users’ ID-based feature representations[9], and such recommender systems perform poorly in the cold-start problem of models. Therefore, semantics can be utilized to enhance the cold start of the model. For example, two users can establish an association by the similarity between their textual features.
In this paper, we propose a novel framework for personalized federated recommendation, termed UFGraphFR (User Text description-based Feature Graph Federated Recommendation), which constructs a user relationship graph based on structured textual features without exposing private data. The proposed framework enables the server to aggregate user embeddings and generate personalized recommendations while complying with the privacy-preserving principle of federated learning. Specifically, our main contributions are summarized as follows:
  • 1.
    To better utilize user-structured attribute data, we introduce a joint-embedding layer that transforms user-side structured features into natural language descriptions, encodes them via a pre-trained language model into high-dimensional embeddings, and further projects them into compact representations via a lightweight linear transformation. The projection weights are shared with the server to guide semantic graph formation.
  • 2.
    To address the limitation that existing federated recommender systems ignore correlations between users, we propose a user graph construction method that builds a semantic user relationship graph on the server using differentially private embedding weights of user text descriptions. This allows the system to model collaborative preferences while maintaining strict data privacy.
  • 3.
    To validate the effectiveness of the proposed framework, we conduct extensive experiments on multiple real-world datasets. Results show that UFGraphFR achieves superior performance compared to existing federated and centralized baselines while preserving user privacy. And verify the effectiveness under different pre-trained language model embeddings.
@@ 2.1 Federated Recommendation Systems
Federated Recommendation Systems (FedRS, as shown in Fig.1) is an important extension of Federated Learning (FL) in the field of recommendation, which realizes cross-client personalization modeling by protecting the user’s privacy. Personalized modeling, allows the user’s raw data to complete the training process of the recommender system without leaving the client while protecting the user’s privacy.
Some existing basic recommender system approaches (e.g., Matrix-Factorization-Based Recommender Systems[10], Neural-Collaborative-Filtering-Based Recommender Systems[11]) have been combined with the FL framework to form, for example, matrix factorization-based federated recommender systems FCF[12], FedMF[4], MetaMF[13], FedRecon[14], and neural collaborative filtering-based federated recommender system FedNCF[15].After this, more and more FedRSs were proposed. Zhang et al.[16] proposed a personalized FedRec framework, PFedRec, which removes user embedding and learns personalized scoring functions to capture user preferences. However, this framework ignores the correlation between users and does not respond well to the common preferences among similar users. To compensate for this shortcoming of correlation between users, Liu et al.[17] enhanced the federated recommender system by introducing the user’s social network. However, this framework exposes the user’s social network explicitly, which goes against the premise of protecting user privacy. Wu et al.[18] proposed FedPerGNN, an approach to capture user-item correlations using local map neural networks. At the same time, FedPerGNN also uses servers to find higher-order neighbors, thus providing more favorable information for local model training. However, this approach aligns historical interactions with the server leads to high computational overhead, and increases the risk of exposing users’ private information.
@@ 4.1 Local Training
Figure 2: Client Model:The image shows our client model, with the yellow box on the left indicating the structure of the joint embedding layer (consisting of the prompt layer, pre-trained language model layer, and linear layer used to convert structured attributes into natural language). On the right is the Transformer Block used for sequence modeling
@@ 4.1.1 Joint Embedding Layer
Prompts Output Layer: Each user $u$, has a structured data $P_{u}$, as shown in Fig. 3. The role of Prompts Output Layer is to transform the structured data into textual language.
Figure 3: Prompts Output Layer: Structured data of the user on the left, textual description on the right
@@ 4.1.2 Transformer Block
These embeddings are concatenated into a sequence matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$, which serves as the input to the Transformer Block.
Multi-Head Self-Attention Mechanism To capture the complex dependencies within user interaction sequences, we adopt a multi-head self-attention mechanism. Specifically, for each item embedding in the sequence, we compute contextualized representations by projecting the input matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$ into query ($Q$), key ($K$), and value ($V$) spaces through trainable linear transformations:
@@ 4.2 Federated Training
Figure 4: Service Model: First, it receives the parameters passed up from the client; second, it constructs the user relationship graph based on the parameters of the Jointed-Embedding layer; third, it learns the global item embedding; finally, it sends the global item embedding down to the client.
@@ 4.2.2 Learning global item embedding on the server
Where A is the round item embedding matrix, the I-th row represents the item embedding obtained from user i, $\mathcal{R}$ is the learned correlation matrix, and the i-th behavior $r_{i}$ In addition, l represents the number of convolution layers, l=1 for simple calculations in this model.
@@ 4.4 Algorithm
Algorithm 1 Graph Federation Recommendation System based on User Text description features.
  Initialize $\eta,\lambda,\{\theta_{item,i},\theta_{user,i},\theta_{score,i},\theta_{Transformer,i}\}_{i=1}^{N}$
  for each round $t=1,2,...,T$ do
        Client local training:
            for each client $i=1,2,...,N$ in users do:
                for each local epoch e from 1 to $E$ do:
                    Update($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$) with Eq.(18)
                end for
                    ($\theta^{(t+1)}_{item,i},\theta^{(t+1)}_{user,i},\theta^{(t+1)}_{score,i},\theta^{(t+1)}_{Transformer,i}$) $\leftarrow$ ($\theta^{(t)}_{item,i},\theta^{(t)}_{user,i},\theta^{(t)}_{score,i},\theta^{(t)}_{Transformer,i}$)
            end for
        Server update with graph:
                Expand the user joint-embedding layer weight matrix into a vector
                Calculate the matrix of similarity of user text feature descriptions with Eq.(15)
                Build user relationship graph.
                Learn user common item embeddings with Eq.(16)
                Learn globally shared item embedding $\theta_{global}$ with Eq.(17)
@@ 4.5 Privacy Protection
Where, $\alpha$ is the noise intensity, the larger the $\alpha$, the more noise it introduces, the better the privacy protection performance.
@@ 5.1 Datasets and Evaluation Protocols
We validated the proposed UFGraphFR on four recommended benchmark datasets: MovieLens-100K, MovieLens-1M[41], Lastfm-2K[42] , HetRec2011[42] and Douban [43]. In particular, two MovieLens datasets were collected from the MovieLens website, recording user ratings for movies, with no fewer than 20 ratings per user. Lastfm-2K is a music dataset where each user keeps a list of artists listened to and a listen count. We removed users with less than 5 interactions from Lastfm-2K. HetRec2011 is an extension of MovieLens-10M, which connects movies with the corresponding web pages of the Internet Movie Database (IMDb) and the Rotten Tomatoes movie review system. For the Lastfm-2K and HetRec2011 datasets, we used the user id and the total number of user interaction items as user attributes. Detail statistics are shown in Table.1
@@ 5.2.1 Baselines
  • 1.
    Matrix Factorization (MF)[46] A typical recommendation model. It decomposes the scoring matrix into two embeddings in the same hidden space, describing user and item characteristics respectively.
  • 2.
    Neural Collaborative Filtering (NCF)[47] This method is one of the most representative neural recommendation models. We first learn the user embedded module and the item embedded module, and then model the user’s item interaction by MLP.
  • 3.
@@
  • 4.
    FedMF[4] FedMF trains users locally for embedding and uploads project gradients to the server for global aggregation.
  • 5.
@@
  • 9.
    Federated LightGCN (FedLightGCN)[1] The way LightGCN[7] extends to the federated learning framework in paper GPFedRec. In particular, each client trains local LightGCN using a first-order interaction subgraph.
  • 10.
    Federated Graph Neural Network (FedPerGNN)[18] It deploies a graph neural network on each client, and users can integrate high-level user-project information through a graph extension protocol.
  • 11.

### v4

@@ Abstract
Federated learning offers a privacy-preserving framework for recommendation systems by enabling local data processing; however, data localization introduces substantial obstacles. Traditional federated recommendation approaches treat each user as an isolated entity, failing to construct global user relationship graphs that capture collaborative signals, which limits the accuracy of recommendations. To address this limitation, we derive insight from the insight that semantic similarity reflects preference. similarity, which can be used to improve the construction of user relationship graphs. This paper proposes UFGraphFR, a novel framework with three key components: 1) On the client side, private structured data is first transformed into text descriptions. These descriptions are then encoded into semantic vectors using pre-trained models; 2) On the server side, user relationship graphs are securely reconstructed using aggregated model weights without accessing raw data, followed by information propagation through lightweight graph neural networks; 3) On the client side, user behavior sequences are personalized using Transformer architectures. Extensive experiments conducted on four benchmark datasets demonstrate that UFGraphFR significantly outperforms state-of-the-art baselines in both recommendation accuracy and personalization. The framework also maintains robustness across different pre-trained models, as evidenced by the consistent performance metrics obtained. This work provides a practical method for efficient federated recommendations with strict privacy by using semantic vectors, secure user relationship graphs, and personalized behavior sequences. The code is available at: https://github.com/trueWangSyutung/UFGraphFR.
@@
Federated Learning, Federated Recommendation System, Recommendation
@@ 1 Introduction
In the era of big data and privacy computing, intelligent recommendation systems have become an indispensable infrastructure for modern digital services, enabling personalized movie suggestions, targeted food delivery recommendations, and even emotion-aware content curation through user comment analysis[1]. However, conventional centralized paradigms that aggregate raw user data on servers pose critical privacy risks, as evidenced by growing user reluctance to share personal information with service providers[2]. Federated learning emerges as a promising privacy-preserving framework, allowing collaborative model training without exposing users’ private data [3]. Its application in recommender systems enables the construction of complete models while keeping raw data localized on client devices[4].
However, a fundamental limitation of existing federated recommendation approaches stems from the isolation of raw user-item interaction records on local devices, which prevents the server from accessing the global user-item interaction matrix[5, 6]. Consequently, unlike traditional graph neural network recommendation models that leverage global graph structures (such as user-item bipartite graphs or user social graphs) to capture complex user correlations and collaborative signals, federated systems primarily rely on limited local data from individual users. This constraint inherently caps the upper bound of recommendation accuracy, as models cannot effectively exploit the rich relational patterns available in centralized settings.
In this paper, we propose a novel framework for personalized federated recommendation, termed UFGraphFR (User text-description Feature Graph Federated Rec), which addresses a key challenge: how to find a ’proxy’ signal that enables the server to measure user similarity and approximately build a ’user relationship graph’ without exposing raw private data. To solve this problem, UFGraphFR securely transforms local private structured attributes into textual descriptions and maps them into a unified semantic space using pre-trained language models, thereby enabling safe user relationship construction. The main contributions of this work are summarized as follows:
  • We innovatively leverage user text descriptions as a secure bridge for constructing user relation graphs, breaking away from ID-centric paradigms and enabling semantic-aware similarity modeling under privacy constraints.
  • We design a learning architecture: on the client side, private structured data is first converted into text descriptions and encoded into semantic vectors using pretrained models; on the server side, user relation graphs are reconstructed securely using aggregated model weights without accessing raw data, followed by information propagation via lightweight graph neural networks.
  • We propose the complete UFGraphFR framework, which further incorporates Transformer architectures on the client side to capture personalized user behavior sequences.
  • Through comprehensive experiments on four benchmark datasets, we validate the superiority of UFGraphFR in recommendation accuracy, personalization capability, and robustness, supported by in-depth ablation studies.
The rest of this paper is organized as follows: Section II reviews the related work on federated recommender systems and graph-based learning recommender systems. Section III introduces the UFGraphFR method proposed in this paper, including its overall framework and key technical details. Section IV will focus on the experimental part, including performance experiments and ablation experiments.
@@ 2 Related Work
Table 1: Comparison of related work.
Method Core Idea Pros Cons/Limitations
FedMF [4] Applies matrix factorization within the FL framework. Users train locally and upload gradients. Simple and efficient; Establishes a baseline for FL in recommendation. Ignores user-user correlations; Limited by the linear nature of MF.
FedNCF [7] Extends Neural Collaborative Filtering to FL, using MLPs for non-linear user-item interaction modeling. Captures non-linear interactions; More expressive than linear models like FedMF. Still treats users in isolation; High communication cost due to large MLP parameters.
PFedRec [8] Removes user embeddings from the server, learning personalized scoring functions on each client. Achieves high personalization; Decouples user and item learning. Fails to model collaborative signals between similar users; Performance may degrade for users with sparse data.
FedRecon [9] Retrains user embeddings in every communication round based on local data. Enhances personalization by avoiding stale embeddings. Computationally expensive for clients; May lead to unstable convergence.
FedPerGNN [10] Deploys GNNs on clients and uses a graph extension protocol to find higher-order neighbors. Leverages structural information for better recommendations. Requires aligning historical interactions with the server, risking privacy leaks; High computational and communication overhead.
GPFedRec [1] Proposes a graph-guided aggregation mechanism to capture user preference correlations. Explicitly models user-user relationships; Improves performance by leveraging neighborhood information. The method for constructing the user graph is not explicitly detailed in the provided text, potentially relying on sensitive data.
EasyRec [11] A centralized system using a text-behavior alignment framework with contrastive learning and language model tuning. Effectively integrates rich semantic information from text; Strong generalization capability. Centralized architecture violates user privacy; Not applicable to privacy-sensitive federated settings.
  • This table highlights the trade-offs between personalization, privacy, and the ability to model user correlations in current approaches, motivating the need for a method like UFGraphFR that addresses these limitations
@@ 2.1 Federated Recommendation Systems
Federated Recommendation Systems (FedRS, as shown in Fig.1) is an important extension of Federated Learning (FL) in the field of recommendation, which realizes cross-client personalization modeling by protecting the user’s privacy. Personalized modeling, allows the user’s raw data to complete the training process of the Recommender system(RecSys) without leaving the client while protecting the user’s privacy.
Some existing basic recommender system approaches (e.g., Matrix-Factorization-Based RecSys[12], Neural Collaborative Filtering Based RecSys[13]) have been combined with the FL framework to form, for example, matrix factorization based federated recommender systems FCF[14], FedMF[4], MetaMF[15], FedRecon[9], and neural collaborative filtering based federated recommender system FedNCF[7]. After this, more and more FedRSs were proposed. Zhang et al.[8] proposed a personalized FedRec framework, PFedRec, which removes user embedding and learns personalized scoring functions to capture user preferences. However, this framework ignores the correlation between users and does not respond well to the common preferences among similar users. To compensate for this shortcoming of correlation between users, Liu et al.[16] enhanced the federated recommender system by introducing the user’s social network. However, this framework exposes the user’s social network explicitly, which goes against the premise of protecting user privacy. Wu et al.[10] proposed FedPerGNN, an approach to capture user-item correlations using local map neural networks. At the same time, FedPerGNN also uses servers to find higher-order neighbors, thus providing more favorable information for local model training. However, this approach aligns historical interactions with the server leads to high computational overhead, and increases the risk of exposing users’ private information.
@@ 2.2 Graph Learning-based Recommendation System
To provide a systematic overview of the existing federated recommendation landscape, Table 1 summarizes and compares the key characteristics, advantages, and limitations of several representative methods discussed in this section.
@@ 3 Preliminary
Table 2: Summary of key notations.
Symbol Definition
$U,N$ Set and number of users.
II Set of items.
$r_{ui}$ Interaction between user $u$ and item $i$.
$f(\theta)$ Recommendation model with parameters $\theta$.
$\theta_{\text{user}},\theta_{\text{item}}$ User and item embedding parameters.
$\theta_{\text{global}},\theta_{\text{local}}$ Global shared and local private model parameters.
$\mathcal{G(U,E)}$ User relationship graph with adjacency matrix $\mathcal{A}$.
$\mathcal{L}_{i}(\theta)$ Local loss on client $i$.
$\omega_{i}$ Client $i$’s aggregation weight.
$P_{u},v_{u}$ User $u$’s attribute and PLM-generated text embedding.
$e_{u}$ Final low-dimensional user embedding.
$d,d_{1},d_{2}$ Dimensions: joint embedding, PLM output, linear layer output.
$L$ Length of user’s interaction sequence.
$e_{i_{t}}$ Embedding of the $t$-th interacted item.
$d_{\text{item}}$ Item embedding dimension.
$K,d_{h}$ Number and size of Transformer attention heads.
$\eta,\lambda$ Learning rate and regularization coefficient.
$\alpha$ Noise scale for differential privacy.
@@
For clarity, we summarize the key notations used throughout this paper in Table 2. These symbols are fundamental to the mathematical formulation and description of the proposed UFGraphFR framework.
@@ 4 Methodology
In the following sections, we will provide a detailed introduction to the proposed UFGraphFR, as illustrated in Figure 2.
Figure 2: This figure shows a recommendation system framework that integrates federated learning and graph neural networks. Local training of users and item embeddings on the client side, uploading model parameters to the server; The server constructs a relationship graph based on user similarity, aggregates and generates global item embeddings, and sends them back to the client to achieve collaborative recommendation under privacy protection.
@@ 4.1.1 Joint Embedding Layer
Prompts Output: Each user $u$, has a structured data $P_{u}$. The role of Prompts Output is to transform the structured data into textual language.
@@
Figure 3: Client Model:The image shows our client model, with the yellow box on the left indicating the structure of the joint embedding layer (consisting of the pre-trained language model layer, and linear layer). On the right is the Transformer Block used for sequence modeling.
@@ 4.1.2 Transformer Block
[…]
Multi-Head Self-Attention Mechanism To capture the complex dependencies within user interaction sequences, we adopt a multi-head self-attention mechanism. Specifically, for each element embeddeded in the sequence, we compute contextualized representations by projecting the input matrix $X\in\mathbb{R}^{L\times d_{\text{item}}}$ into theto the query ($Q$),KK)($K$) and value ($V$) spaces through trainable linear transformations:
@@ 4.2.2 Learning global item embedding on the server
where A is the round item embedding matrix, the I-th row represents the item embedding obtained from user i, $\mathcal{R}$ is the learned correlation matrix, and the i-th behavior $r_{i}$ . In addition, l represents the number of convolution layers, l=1 for simple calculations in this model.
@@ 4.5 Privacy Protection
where $\alpha$ is the noise intensity, the larger the $\alpha$, the more noise it introduces, the better the privacy protection performance.
@@ 5 Experiment
Algorithm 1 The Proposed UFGraphFR Algorithm
0: Number of clients NN, total communication rounds $T$, local epochs $E$, learning rate $\eta$, regularization coefficient $\lambda$, user joint-embedding layer parameters $\theta_{user}$, item embedding layer parameters $\theta_{item}$, and other local model parameters $\theta_{local}$.
0: Optimized global item embedding $\theta_{global}$.
1:Initialize global parameters $\theta_{global}^{(0)}$ (e.g., item embeddings) on the server.
2:Broadcast $\theta_{global}^{(0)}$ to all NN clients.
3:for each communication round t=1,2,,Tt=1,2,...,T do
4:  Client Local Training (Parallel Execution):
5:  for each client i=1,2,,Ni=1,2,...,N do
6:   Initialize local item embedding $\theta_{item,i}^{(t)}\leftarrow\theta_{global}^{(t-1)}$.
7:   Use local data to train the local model for $E$ epochs by minimizing the loss function $\mathcal{L}_{all}$ (Eq. (12)) via SGD (Eq. (18)).
8:   Obtain updated local parameters: $\{\theta_{item,i}^{(t)},\theta_{user,i}^{(t)},\theta_{Transformer,i}^{(t)},\theta_{umlp,i}^{(t)},\theta_{score,i}^{(t)}\}$.
9:   Upload the user joint-embedding weight matrix $W_{i}^{(t)}\in\mathbb{R}^{d_{1}\times d_{2}}$ and the local item embedding $\theta_{item,i}^{(t)}$ to the server.
10:  end for
11:  Server Update with Graph Aggregation:
12:  Vectorize: Flatten each client’s uploaded joint-embedding weights: $w_{i}^{(t)}=\text{vec}(W_{i}^{(t)})\in\mathbb{R}^{d_{1}d_{2}}$.
13:  Compute Similarity: Calculate cosine similarities $\mathcal{S}_{ij}^{(t)}$ between all $w_{i}^{(t)}$ and $w_{j}^{(t)}$.
14:  Build Graph: Construct user graph $\mathcal{G}^{(t)}$ by connecting each user to its top-$k$ similar users. Obtain adjacency matrix $\mathcal{A}^{(t)}$ and degree matrix $\mathcal{D}^{(t)}$.
15:  Aggregate Embeddings: Apply GCN to local item embeddings $A^{(t)}$: $\mathcal{R}^{(t)}=(\mathcal{D}^{(t)})^{-1/2}\mathcal{A}^{(t)}(\mathcal{D}^{(t)})^{-1/2}A^{(t)}$.
16:  Update Global Embedding: Set $\theta_{\text{global}}^{(t)}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{R}_{i}^{(t)}$.
17:  6. Broadcast $\theta_{\text{global}}^{(t)}$ to all clients.
18:end for
19:return The final global item embedding $\theta_{global}^{(T)}$.
@@ 5.1 Datasets and Evaluation Protocols
We validated the proposed UFGraphFR on four recommended benchmark datasets: MovieLens-100K, MovieLens-1M[37], Lastfm-2K[38] and HetRec2011[38]. In particular, two MovieLens datasets were collected from the MovieLens website, recording user ratings for movies, with no fewer than 20 ratings per user. Lastfm-2K is a music dataset where each user keeps a list of artists listened to and a listen count. We removed users with less than 5 interactions from Lastfm-2K. HetRec2011 is an extension of MovieLens-10M, which connects movies with the corresponding web pages of the Internet Movie Database (IMDb) and the Rotten Tomatoes movie review system. For the Lastfm-2K and HetRec2011 datasets, we used the user id and the total number of user interaction items as user attributes. Detail statistics are shown in Table.3
@@ 5.2.1 Baselines
  • Matrix Factorization (MF)[41] A typical recommendation model. It decomposes the scoring matrix into two embeddings in the same hidden space, describing user and item characteristics, respectively.
  • Neural Collaborative Filtering (NCF)[42] This method is one of the most representative neural recommendation models. We first learn the user-embedded module and the item-embedded module, and then model the user’s item interaction by MLP.
@@
  • FedMF[4] FedMF trains users locally to embed and uploads project gradients to the server for global aggregation.
@@
  • Federated LightGCN (FedLightGCN)[1] The way LightGCN[5] extends to the federated learning framework in the GPFedRec paper. In particular, each client trains local LightGCN using a first-order interaction subgraph.
  • Federated Graph Neural Network (FedPerGNN) [10] It deploys a graph neural network on each client, and users can integrate high-level user-project information through a graph extension protocol.